According to the ideal gas law, a 10.74 mol sample of krypton gas in a 0.8444 L container at 498.7 K should exert a pressure of 520.5 atm. What is the percent difference between the pressure calculated using the van der Waals' equation and the ideal pressure? For Kr gas, a = 2.318 L2atm/mol2 and b = 3.978×10-2 L/mol. ____%
Answer #1

Van der waals equation is given by,

P = RT/(Vm-b) - a/Vm^2

Vm = V/n = 0.8444/10.74 = 0.07862

P = 0.0821*498.7/0.07862 -0.03978. - 2.318/(0.07862)^2

P = (40.94327/0.03884) - 375.01

P = 1054.15 -375.01 = 679.14atm

Percentage difference = 679.14 - 520.5 / 520.5 * 100 = 30.5%